OpenStudy (anonymous):

What is the sum of the geometric sequence 1, 3, 9, … if there are 10 terms

OpenStudy (raden):

the first term a1 = 1 with ratio r = a2/a1 = 3/1 = 3 to get the sum of that series use the formula : S10 = a1(r^n - 1)/(r - 1)

OpenStudy (raden):

i mean Sn = a1(r^n - 1)/(r - 1) with n = 10

OpenStudy (anonymous):

Do you know the formula for a geometric progression? It's \[ar ^{n-1}\] Here, a = the first in the series = 1 r = the ratio = 3/1 = 9/3 = 3 n = 10 Now plug in the values.

OpenStudy (anonymous):

These are my choices 29,524 55,987 87,381 88,573

OpenStudy (raden):

therefore, calculate this S10 = 1(3^10 - 1)/(3 - 1) = ...

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

I got 59047.7 ????

OpenStudy (anonymous):

I suck at math so much!

OpenStudy (anonymous):

Wait its 29.524!

OpenStudy (raden):

finally, u got it :) yes, it is 29,524

OpenStudy (anonymous):

Thank you

OpenStudy (raden):

you're welcome